The Good Seed Flower Shop sells its own brand of fertilizer. It mixes one brand of fertilizer that contains 7% nitrogen with another than contains 8% nitrogen. The resulting mixture contains 7.7% nitrogen. How many pounds of the 7% fertilizer must be used to make 400 pounds of the 7.7% mixture?
Answer by mananth(16946)

The Good Seed Flower Shop sells its own brand of fertilizer. It mixes one brand of fertilizer that contains 7% nitrogen with another than contains 8% nitrogen. The resulting mixture contains 7.7% nitrogen. How many pounds of the 7% fertilizer must be used to make 400 pounds of the 7.7% mixture
7%-------------------- 8%----------------mix 7.7%
x----------------------400-x-------------- 400 ----------- lbs

0.7x+0.8(400-x)=400*0.77
0.7x+320-0.8x=308
-0.1x=-12
x=120
120--------------------380-----------------400 Answer